Sol Summary – May 3rd

Crew 261 Sol Summary Report 03-05-2023

Sol: 3

Summary Title: experiencing EVA procedures in the windy environment of Mars

Author’s name: Aline Decadi, XO

Mission Status: Nominal.

Sol Activity Summary: Today took place 2 EVAs:
– EVA-3 with the following crew members: James Burk (Commander) (EVA Leader), Cecile Renaud (GreenHab Officer).
– EVA-4 with the following crew members: Aline Decadi (XO) (EVA Leader), Erin Kennedy (Crew Roboticist).

EVA-3 conducted a collection of soil for GreenHab experiments and operational testing of EVALink and VR-Comms procedures using Garmin devices. The soil collection is meant to support our Biostimulation experiments.

The crew drove with the Opportunity rover towards Cow Dung Road, took it South to the entry to Robert’s Rock Garden, and walked East to the collection site.

EVA-4 conducted Atmosphinder testing using the half-sized robot (~1 m diameter) in front of the Hab. The objectives of the test focused on
structural aspects of the robot and de-risking operational elements. The EVA crew members worked together to increase the operational understanding of extracting the robot out of the RAM, loading the robot into the rover (ATV), securing the robot to the rover (ATV), and driving slowly with sudden decelerations. The lessons learned from this exercise will de-risk these elements on future EVAs when transporting the robot to regions of interest. Quantitative testing of the robot was completed using force meters to measure the longitudinal and latitudinal forces required to lift the robot. Qualitative testing observed that the robot can roll down a small incline unassisted.
Structural elements on the robot that failed were known previously and can be repaired. The electronics payload was tested separately, with the new addition of the wind sensor compared to EVA-2. The EVA-4 worked smoothly, and valuable new information was learned. The success of this EVA-4 was in large part thanks to the EVA crew lead and the crew facilitating behind the scenes.

We spent the afternoon making analysis and processing yesterday’s experiments: the Musk observatory first Sun imaging; analyzing the soil samples collected for the GreenHab and Biostimulation experiments; and setting up the tomorrow’s experiments: robotics Atmosphinder, drone reconnaissance of campus with 4 waypoints. The Commander James Burk also made a detailed demonstration on MarsVR, showing the MDRS and the surrounding desert with very accurate details of the terrain, and also how we will be able to use the VR to plan out our EVAs and even see EVAs in progress.

The cardiovascular monitoring with requested parameters has been also performed as usual by HSO.

Biostimulation experiments are divided in 2 sections:
– early growth biostimulation using different kinds of soil (Utah soil, Martian
regolith simulant and normal soil).
– biostimulation on grown tomato plants.
Biostimulant solution is made from spirulina.

We spent the evening having dinner, the commander briefing, and preparing the experiments for tomorrow.

Look Ahead Plan: We will be in Sim tomorrow to perform a flyover of the campus using a drone, and to test out the Meshtastic devices for the VR, and make a complementary structural test for the Atmosphinder experiment, according to mission request approval.

Anomalies in work: Toilet is smelly, and Crew Engineer will cover that in Operations report.

Weather: Sunny and pleasant. Very windy day.

Crew Physical Status: Nominal

EVA: testing Atmosphinder prototype, drone reconnaissance of campus.

Reports to be filed: HSO Report, EVA Request, EVA Report, Operations Report, Journalist Report, Pictures of the day.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site